1. A station can process one unit every 5 minutes, but the station must be taken down for maintenance for the last 10 minutes of every hour.   What is the effective production rate at the station, in units per hour?   In the answer box, simply include the number (i.e. you do not need to write units/hour).
Answer #1

Effective production rate is the rate of units per hour that can be processed or produced at a station given the constraints like maintainance etc.

Given :

1 unit is processed in 5 minutes. In 1 minute 1/5 units are processed.

We have 50 minutes of effective working time at the station (per hour)

Effective production rate = effective production time per hour(in minutes) * units processed per minute = 50 * 1/5=10

10 units is the effective production rate.

10.
Answer #2

To calculate the effective production rate at the station, we need to take into account the downtime for maintenance.

The station can process one unit every 5 minutes, which means it can process 12 units in an hour (since there are 60 minutes in an hour and 60/5 = 12). However, it must be taken down for maintenance for the last 10 minutes of every hour.

So, in one hour, the station can only operate for 60 minutes - 10 minutes = 50 minutes.

Now, we can calculate the effective production rate:

Effective production rate = Number of units processed per hour / Operating time per hour Effective production rate = 12 units / 50 minutes

To express the answer in units per hour, we need to convert the time to hours:

Effective production rate = 12 units / (50 minutes / 60 minutes per hour) Effective production rate = 12 units / (5/6) hours Effective production rate = 12 units * (6/5) hours Effective production rate = 14.4 units per hour

So, the effective production rate at the station is 14.4 units per hour.

answered by: Hydra Master